# Profiling stellar environments of gravitational wave sources

### v1

@@ Abstract
Discussion.— The astrophysical processes that govern compact binary coalescences (CBCs) are at best partially understood, if at all Spera et al. (2022). Several models of the environments in which CBCs form, evolve, and merge have been proposed Mapelli (2021), although, there is no conclusive evidence to suggest that one formation channel is preferred over others. Indeed, some works suggest that multiple formation channels have contributed to the production of the detected stellar-mass CBCs to date Zevin et al. (2021); Pierra et al. (2024). The intrinsic parameters of a CBC, viz., its component masses, spins, and/or orbital eccentricity, can sometimes be suggestive of a subset of possible merger environments and formation channels Farr et al. (2017); Bavera et al. (2020); Bogdanovic et al. (2007); Vajpeyi et al. (2022); McKernan and Ford (2024); Adamcewicz et al. (2024); Samsing et al. (2022); Tagawa et al. (2021); Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Yu et al. (2020); Randall and Xianyu (2018); Gerosa and Fishbach (2021); Liu and Lai (2021). They cannot, however, always conclusively rule out all other possible formation channels, especially if there is no EM counterpart.
@@
In this work we show for the first time that, on a single-event basis, rich and precise information of the environment with generic gravitational potentials can be extracted. Additionally, we show that it also possible to acquire information about the location of the CBC with respect to the center of the potential. This can be achieved provided the CBC’s LOSA and some of its higher time derivatives can be measured. We consider the SMBH, NSC (BW), and GC (Plummer) potentials. We find that we can measure enclosed mass and potential profile exponents, and the eccentricity and the radius of the CBC’s orbit in the potential, with relative errors of less than unity for a large portion of the parameter space we consider, even going down to sub-percent precision for certain portions.
@@
We expect that our work, demonstrating the possibility of constraining environment parameters, on a single-event basis, could have significant astrophysical implications, potentially answering the longstanding question of the provenance of CBCs for at least a fraction of them with measurable LOSV time derivatives [^11] See Tiwari et al. (2024b) for estimates of this fraction (focusing exclusively on LOSAs) for various GC configurations.. Moreover, it could unearth important correlations between the mass of the CBC, the mass of the environment, the steepness of the potential profile, and the location of the CBC with respect to the center of the potential. Probing these would directly test astrophysical mechanisms, such as dynamical friction and mass segregation, among others, in dense stellar environments.
@@
Strictly, our work probes the environment near the end of the CBC’s lifetime. Given that the delay time between the CBC’s formation and merger could span millions to billions of years (see, e.g., Fishbach and Kalogera (2021), and references therein), the location of the environment probed by the CBC could be quite different from the location of its formation, due to kicks induced by, e.g., asymmetric supernova mechanisms, or interactions with other bodies [^12] The probability of such interactions are relatively higher in dense star clusters such as NSCs.. Thus, our work complemented by EM observations of stellar binaries could enable the construction of the complete story of a binary’s evolution, starting out as well-separated (bound) main-sequence stars and ending with the coalescence of compact objects.
@@ Acknowledgements.
We thank Debatri Chattopadhyay for their careful reading and feedback. We thank K. G. Arun, Debarati Chatterjee and Zoheyr Doctor for useful discussions. SJK gratefully acknowledges support from SERB Grants SRG/2023/000419 and MTR/2023/000086. SC acknowledges support from the Department of Atomic Energy, Government of India, under project no. 12-R&D-TFR-5.02-0200 and RTI 4002. All computations were performed on the Sarathi Computing Cluster at IUCAA.

### v2

@@ IV Discussion
The astrophysical processes that govern compact binary coalescences (CBCs) are at best partially understood, if at all Spera et al. (2022). Several models of the environments in which CBCs form, evolve, and merge have been proposed Mapelli (2021), although there is no conclusive evidence to suggest that one formation channel is preferred over others. Indeed, some works suggest that multiple formation channels have contributed to the production of the detected stellar-mass CBCs to date Zevin et al. (2021); Pierra et al. (2024). The intrinsic parameters of a CBC, viz., its component masses, spins, and/or orbital eccentricity, can sometimes be suggestive of a subset of possible merger environments and formation channels Farr et al. (2017); Bavera et al. (2020); Bogdanovic et al. (2007); Vajpeyi et al. (2022); McKernan and Ford (2024); Adamcewicz et al. (2024); Samsing et al. (2022); Tagawa et al. (2021); Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Yu et al. (2020); Randall and Xianyu (2018); Gerosa and Fishbach (2021); Liu and Lai (2021). They cannot, however, always conclusively rule out all other possible formation channels, especially if there is no EM counterpart.
@@
In this work, we show that rich and precise information of the environment with generic gravitational potentials can be extracted on a single-event basis. Additionally, we show that it is also possible to acquire information about the location of the CBC with respect to the center of the potential. This can be achieved provided the CBC’s LOSA and some of its higher time derivatives can be measured. We consider the SMBH, NSC (BW), and GC (Plummer) potentials. We find that we can measure enclosed mass and potential profile exponents, and the eccentricity and the radius of the CBC’s orbit in the potential, with relative errors of less than unity for a large portion of the parameter space we consider, even going down to sub-percent precision for certain portions.
An important assumption inherent in our analysis is that we know the true model (environment + outer-orbit) to fit to the data. This, of course, is not the case in general. Moreover, as we show in Figure 11 of the Appendix D, fitting a wrong model to the data will result in potential stealth biases. Nevertheless, such biases can be avoided by selecting the correct model via Bayesian model selection. By evaluating the Savage-Dickey density ratio Zimmerman et al. (2019), we demonstrate in Figure 7 that it should be possible to differentiate the true orbit from other incorrect alternatives, given the same potential. We also perform a full Bayes factor analysis and show that two different potentials can also be confidently distinguished.
@@
We expect that our work, demonstrating the possibility of constraining environment parameters, on a single-event basis, could have significant astrophysical implications, potentially answering the longstanding question of the provenance of CBCs for at least a fraction of them with measurable LOSV time derivatives (see Tiwari et al. (2024) for estimates of this fraction , focusing exclusively on LOSAs, for various GC configurations). Moreover, it could unearth important correlations between the mass of the CBC, the mass of the environment, the steepness of the potential profile, and the location of the CBC with respect to the center of the potential. Probing these would directly test astrophysical mechanisms, such as dynamical friction and mass segregation, among others, in dense stellar environments.
@@
Strictly, our work probes the environment near the end of the CBC’s lifetime. Given that the delay time between the CBC’s formation and merger could span millions to billions of years (see, e.g., Fishbach and Kalogera (2021), and references therein), the location of the environment probed by the CBC could be quite different from the location of its formation. This is due to kicks induced by, e.g., asymmetric supernova mechanisms, or interactions with other bodies (the probability of which is relatively higher in dense star clusters Spitzer (1987); Binney and Tremaine (2008)).
Thus, our work, complemented by observations of binary stars, could enable the construction of the complete story of their evolution, starting out as well-separated (bound) main-sequence stars, and ending with the coalescence of compact objects.
@@ Acknowledgements.
We thank Debatri Chattopadhyay for their careful reading and feedback. We thank K. G. Arun, Debarati Chatterjee and Zoheyr Doctor for useful discussions. SJK gratefully acknowledges support from SERB Grants SRG/2023/000419 and MTR/2023/000086. SC acknowledges support from the Department of Atomic Energy, Government of India, under project no. 12-R&D-TFR-5.02-0200 and RTI 4002. AV acknowledges support to CITA by the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580). All computations were performed on the Sarathi Computing Cluster at IUCAA.